# DynaSeg: A Deep Dynamic Fusion Method for Unsupervised Image Segmentation Incorporating Feature Similarity and Spatial Continuity

### v3

@@ Abstract
Our work tackles the fundamental challenge of image segmentation in computer vision, which is crucial for diverse applications. While supervised methods demonstrate proficiency, their reliance on extensive pixel-level annotations limits scalability. We introduce DynaSeg, an innovative unsupervised image segmentation approach that overcomes the challenge of balancing feature similarity and spatial continuity without relying on extensive hyperparameter tuning. Unlike traditional methods, DynaSeg employs a dynamic weighting scheme that automates parameter tuning, adapts flexibly to image characteristics, and facilitates easy integration with other segmentation networks. By incorporating a Silhouette Score Phase, DynaSeg prevents undersegmentation failures where the number of predicted clusters might converge to one. DynaSeg uses CNN-based and pre-trained ResNet feature extraction, making it computationally efficient and more straightforward than other complex models. Experimental results showcase state-of-the-art performance, achieving a 12.2% and 14.12% mIOU improvement over current unsupervised segmentation approaches on COCO-All and COCO-Stuff datasets, respectively. We provide qualitative and quantitative results on five benchmark datasets, demonstrating the efficacy of the proposed approach. Code available at https://github.com/RyersonMultimediaLab/DynaSeg
@@ 2 Related Work
Recent advancements in machine learning have propelled various vision applications, notably in classification [19] and object detection [20]. Image segmentation, a crucial task in computer vision, has seen significant progress, primarily driven by supervised learning techniques. However, the reliance on annotated data poses challenges in scalability and cost. This section reviews existing methods and situates our approach within the broader landscape, emphasizing its strengths and addressing the limitations of previous techniques.
2.1 Supervised Segmentation
The advent of convolutional neural networks (CNNs) and large annotated datasets has revolutionized image segmentation, surpassing traditional algorithms such as active contour models (ACM) [9], k-means [10], and graph-based segmentation (GS) [11]. Traditional methods often struggle with noise susceptibility, intricate boundary handling, and sensitivity to data quality [21, 22, 23]. In contrast, CNN-based techniques like U-Net [24], SegNet [25], and DeepLab [26] offer precise segmentation capabilities.
Among the pioneering approaches, the Fully Convolutional Network (FCN) [27] inspired by the VGG network [28] utilized multi-resolution layers for segmentation. However, FCN results often appeared blurry and smooth, needing more precision. Addressing this, Ronneberger proposed U-Net [24], catering to fewer annotated images with reasonable training times, notably in biomedical applications. Further advancements were made with Mask R-CNN [13], extending Faster R-CNN [29] by predicting object masks alongside bounding boxes, significantly improving instance segmentation performance. Following the success of Mask R-CNN, several novel deep neural architectures have been proposed, including transformers [30] and capsule networks [31], both showing promising results. Despite their effectiveness, these models rely heavily on large annotated datasets, escalating annotation costs and introducing challenges related to missing labels and data quality variations [32, 33, 34].
While supervised segmentation methods achieve excellent results on benchmark datasets, their application in real-world scenarios is costly due to the dependence on extensive annotated data. This limitation has catalyzed the development of unsupervised segmentation models, including our proposed method. Our approach provides a scalable and cost-efficient solution, effectively addressing the need for robust segmentation without the necessity for explicit annotations.
Figure 1: DynaSeg: Dynamic Weighting Scheme for Unsupervised Image Segmentation Framework. This framework involves a forward-backward process that is iterated $T$ times to obtain the final prediction of the cluster labels $c_{n}$.
2.2 Unsupervised Segmentation
Recent advancements in unsupervised segmentation methods offer a cost-effective alternative, aiming to discover objects, boundaries, and materials without explicit ground truth. Among these approaches, clustering-based and CNN-based methods have gained prominence, each offering unique strategies to tackle the challenges of unsupervised segmentation.
2.2.1 Clustering-Based Methods
@@
Foundational unsupervised segmentation methods like K-means [10], Mean Shift [35], and graph-based segmentation (GS) [11] form clusters based on pixel feature similarity but often lack spatial coherence. These methods treat each pixel independently, resulting in fragmented segmentations with disconnected regions, especially images with complex structures and textures. In contrast, our model incorporates both feature similarity and spatial continuity to ensure that similar and contiguous pixels belong to the same cluster. Unlike static clustering methods [10][35] that require the number of clusters to be specified a priori, DynaSeg adjusts the number of clusters during training, enhancing the model’s flexibility and robustness.
@@
One significant clustering-based approach is Deep Image Clustering (DIC) [15], which breaks down the segmentation problem into feature transformation and trainable deep clustering sub-networks. However, it relies on superpixels, introducing predetermined boundaries. Similarly, Invariant Information Clustering (IIC) [16], and its versions [36] uses mutual-information-maximization for clustering. However, its reliance on enforcing a uniform distribution over clusters makes it particularly effective with well-balanced datasets [33, 34]. Yet, the efficacy of IIC [16] is contingent upon data balance and can be influenced by the specific data augmentation strategy employed [37, 38, 39]. Unlike the methods [15][16] that separate feature extraction and clustering into distinct steps, DynaSeg employs Joint Optimization where the backpropagation of clustering losses directly influences the CNN. This integration ensures a cohesive segmentation process that is continuously refined, overcoming the limitations of predetermined boundaries and data imbalance.
2.2.2 Deep Learning-Based Methods
@@
Deep learning-based methods have garnered significant attention in unsupervised image segmentation, showcasing remarkable capabilities across various approaches. For instance, Generative Adversarial Networks (GANs) [40] [41] [42] [43] have emerged as powerful tools for creating segmented images by effectively separating foreground and background. Specifically, the GAN-based approach proposed by Melas-Kyriazi [40] leverages pre-trained GANs like BigBiGAN [41] to extract salient object segmentation from the latent space. Despite their innovation, these methods often rely on some form of class supervision, which inherently limits their purely unsupervised nature. Labels4Free [42] extends Style-GAN2 [43] with a segmentation branch for unsupervised foreground object segmentation, typically focusing on binary clustering. However, these approaches tend to rely on discriminative losses, such as binary cross-entropy loss, for training the discriminator network, which may not directly optimize for segmentation quality, resulting in suboptimal outcomes. In contrast, our model adopts a hybrid approach for multi-class segmentation by integrating a clustering strategy with CNN-based feature extraction.
@@
Wnet [46] combines U-net [24] structures into auto-encoders for segmentation, followed by a post-processing phase for segmentation refinement. While effective, this method is computationally expensive and requires extensive hyperparameter tuning. Infomax [14] presents an alternative where an input image is partitioned into superpixels. The Region-Wise-Embedding (RWE) extracts a feature embedding for each superpixel region, and Mutual-Information-Maximization is employed with adversarial training. Despite its efficiency, it faces challenges like superpixel reliance and predetermined boundaries. In contrast, our proposed hybrid model offers a streamlined solution by eliminating the need for post-processing. Our model generates compact clusters directly from the feature space by dynamically optimizing the clustering loss and spatial continuity loss. This integrated approach enables our model to extract rich and informative features, facilitating accurate and robust segmentation, even in challenging scenarios. DynaSeg offers enhanced flexibility and adaptability by removing the reliance on predefined boundaries or superpixels [47], ensuring more natural and visually appealing segmentation outcomes.
Differentiable feature clustering, as proposed by [17], is a CNN-based algorithm that optimizes pixel labels and feature representations using a combination of feature similarity and spatial continuity constraints [17]. However, achieving the desired segmentation result with this approach often requires manual parameter tuning to find the optimal balancing weight ($\mu$) between feature similarity and spatial continuity. This manual tuning process can be cumbersome and may fail to balance the two constraints well, mainly when dealing with images of varying levels of detail and datasets with different characteristics. Moreover, relying solely on fixed weight factors for feature similarity and spatial continuity constraints can lead to the number of predicted clusters converging to one, making it challenging to adapt to different datasets and requiring extensive fine-tuning. Additionally, the lack of adaptability and the need for manual parameter adjustments make integrating this approach with other networks difficult. In contrast, our model introduces a Dynamic Weighting Scheme that dynamically adjusts the weighting parameter ($\mu$) during iterations. By incorporating a silhouette score-based phase to guide the optimization process, our model prevents under-segmentation failures and ensures robust performance across diverse datasets. This dynamic approach not only simplifies the integration of our model with other networks but also enhances its adaptability and scalability, making it well-suited for various segmentation tasks and datasets.
@@
Additionally, recent advancements in CNN architectures, such as Vision Transformers (ViT) [48], have shown promise in various computer vision tasks, including image segmentation. However, ViTs come with their own set of challenges, such as high computational complexity, data inefficiency, training instability, and a propensity for overfitting, which can limit their accessibility and applicability in specific contexts. In contrast, our hybrid model offers a more efficient and scalable solution by dynamically optimizing the clustering loss and leveraging a CNN-based network or a pre-trained ResNet with FPN to extract high-level features from input images. This approach ensures robustness and adaptability across different datasets. Overall, our model provides a practical and effective solution for unsupervised image segmentation, addressing the limitations of existing methods and presenting promising opportunities for real-world applications.
@@ 3.2.2 Pre-trained ResNet with Feature Pyramid Network (FPN) Decoder
To enhance our network’s feature extraction capabilities, we incorporate a pre-trained ResNet-18 model. This approach addresses the need for more prosperous and diverse image segmentation representations, leveraging the strengths of both ResNet and FPN.
@@ 4.1.2 BSD500
Method Backbone mIoU All
Modified DC [59] - 9.8
IIC [16] ResNet18 6.7
Picie [44] ResNet18 14.4
DenseSiam [45] ResNet18 16.4
DynaSeg - FSF CNN-Based 30.51
DynaSeg - SCF CNN-Based 27.57
DynaSeg - FSF ResNet-18 + FPN 30.07
DynaSeg - SCF ResNet-18 + FPN 30.52
@@ 4.3.3 BSD500 and PASCAL VOC2012 Datasets
Quantitative results on BSD500 and PASCAL VOC2012 datasets are detailed in Table 1. SCF achieves a remarkable mIOU of 0.396 on PASCAL VOC2012, demonstrating its effectiveness. FSF consistently outperforms other methods on BSD500, emphasizing its robustness across various ground truth types. In addition to quantitative evaluations, we provide a demo [62] of incremental segmentation using DynaSeg on a sample image from the Pascal VOC2012 dataset. This demo is available in the supplemental video, showcasing our method’s application to real-world images.
@@ 5 Ablation Study
5.1 Silhouette Score Impact
Complementing the tabular results, the accompanying graph provides insights into the distribution of ground truth cluster numbers. On one side, it contrasts the distribution with a random threshold of 3, revealing a substantial disparity. On the other side, the distribution is compared with the optimal number of clusters derived from the Silhouette Score, demonstrating a strong correlation and underscoring the effectiveness of the proposed dynamic weighting scheme.

### v4

@@ Abstract
Our work tackles the fundamental challenge of image segmentation in computer vision, which is crucial for diverse applications. While supervised methods demonstrate proficiency, their reliance on extensive pixel-level annotations limits scalability. We introduce DynaSeg, an innovative unsupervised image segmentation approach that overcomes the challenge of balancing feature similarity and spatial continuity without relying on extensive hyperparameter tuning. Unlike traditional methods, DynaSeg employs a dynamic weighting scheme that automates parameter tuning, adapts flexibly to image characteristics, and facilitates easy integration with other segmentation networks. By incorporating a Silhouette Score Phase, DynaSeg prevents undersegmentation failures where the number of predicted clusters might converge to one. DynaSeg uses CNN-based and pre-trained ResNet feature extraction, making it computationally efficient and more straightforward than other complex models. Experimental results showcase state-of-the-art performance, achieving a 12.2% and 14.12% mIOU improvement over current unsupervised segmentation approaches on COCO-All and COCO-Stuff datasets, respectively. We provide qualitative and quantitative results on five benchmark datasets, demonstrating the efficacy of the proposed approach. Code available at Code available at https://github.com/RyersonMultimediaLab/DynaSeg
@@ 2 Related Work
Recent advancements in machine learning have significantly impacted various vision applications, particularly image segmentation. The introduction of convolutional neural networks (CNNs) and the availability of large annotated datasets have revolutionized image segmentation, enabling these models to surpass traditional algorithms such as active contour models (ACM) [9], k-means clustering [10], and graph-based segmentation (GS) [11]. Notable CNN-based techniques, including Mask R-CNN [13], U-Net [19], SegNet [20], and DeepLab [21], offer precise segmentation capabilities. However, these models are heavily reliant on large annotated datasets, which not only escalate annotation costs but also introduce challenges related to missing labels and data quality variations [22, 23, 24]. These limitations have driven the development of unsupervised segmentation models. Among these approaches, clustering-based and CNN-based methods have gained prominence, each offering unique strategies to address these challenges. Our model, DynaSeg, exemplifies this trend by providing scalable and cost-efficient solutions without the need for explicit annotations.
2.1 Clustering-Based Methods
@@
Foundational unsupervised segmentation methods like K-means [10], Mean Shift [25], and graph-based segmentation (GS) [11] cluster pixels based on feature similarity but often lack spatial coherence. These methods treat each pixel independently, resulting in fragmented segmentations, especially in images with complex structures. In contrast, our model integrates feature similarity and spatial continuity, ensuring that similar and contiguous pixels form the same cluster. Unlike static methods requiring a predetermined number of clusters [10][25], DynaSeg adjusts the number of clusters during training, enhancing flexibility and robustness.
@@
A notable approach is Deep Image Clustering (DIC) [15], which transforms features and clusters them using deep sub-networks but relies on superpixels with predetermined boundaries. Similarly, Invariant Information Clustering (IIC) [16] and its variants [26] use mutual-information-maximization but require uniform cluster distributions, making them effective with balanced datasets [23, 24]. Their efficacy can be influenced by data augmentation strategies [27, 28, 29]. Unlike methods that separate feature extraction and clustering [15, 16], DynaSeg employs Joint Optimization where the backpropagation of clustering losses directly influences the CNN. This integration ensures a cohesive segmentation process, continuously refining the model and overcoming the limitations of predetermined boundaries and data imbalance.
2.2 Deep Learning-Based Methods
@@
Deep learning-based methods have gained significant attention in unsupervised image segmentation, showcasing various innovative approaches. Generative Adversarial Networks (GANs) [30] [31] [32] [33] have emerged as powerful tools for creating segmented images by separating foreground and background. For example, Melas-Kyriazi [30] leverages pre-trained GANs like BigBiGAN [31] to extract salient object segmentation from the latent space. Despite their innovation, these methods often require some form of class supervision, limiting their purely unsupervised nature. Labels4Free [32] extends Style-GAN2 [33] with a segmentation branch for unsupervised foreground object segmentation, typically focusing on binary clustering. However, these approaches tend to rely on discriminative losses, such as binary cross-entropy loss, for training the discriminator network, which may not directly optimize segmentation quality, resulting in suboptimal outcomes. In contrast, our model adopts a hybrid approach for multi-class segmentation by integrating a clustering strategy with CNN-based feature extraction.
@@
Wnet [36] combines U-net [19] structures into auto-encoders for segmentation, followed by a post-processing phase for refinement. While effective, this method is computationally expensive and requires extensive hyperparameter tuning. Infomax [14] presents an alternative where an input image is partitioned into superpixels. The Region-Wise-Embedding (RWE) extracts a feature embedding for each superpixel region, and Mutual-Information-Maximization is employed with adversarial training. Despite its efficiency, it faces challenges like superpixel reliance and predetermined boundaries. In contrast, our proposed hybrid model offers a streamlined solution by eliminating the need for post-processing. Our model generates compact clusters directly from the feature space by dynamically optimizing the clustering loss and spatial continuity loss. This integrated approach enables our model to extract rich and informative features, facilitating accurate and robust segmentation, even in challenging scenarios. DynaSeg offers enhanced flexibility and adaptability by removing the reliance on predefined boundaries or superpixels [37], ensuring more natural and visually appealing segmentation outcomes.
Figure 1: Overview of the DynaSeg Framework: The Feature Extractor Network generates a feature map, classified into $q$ clusters via a linear classifier and batch normalization, resulting in a normalized response map. Cluster labels $c_{i}$ are assigned to each pixel using the argmax function. The number of clusters $q^{\prime}$ is dynamically updated based on feature similarity and spatial continuity. Loss $L$ is computed during backpropagation, with parameters updated using SGD. This process iterates $T$ times to refine cluster labels $c_{i}$, achieving final segmentation. The Silhouette Score sets $opt\_nC$ as the threshold for $q^{\prime}$ to prevent under-segmentation. Black arrows indicate the feedforward path, while red arrows represent backpropagation.
@@
Additionally, recent advancements in CNN architectures, such as Vision Transformers (ViT) [38], have shown promise in various computer vision tasks, including image segmentation. However, ViTs come with their own set of challenges, such as high computational complexity, data inefficiency, training instability, and a propensity for overfitting, which can limit their accessibility and applicability in specific contexts. In contrast, our hybrid model offers a more efficient and scalable solution by dynamically optimizing the clustering loss and leveraging a CNN-based network or a pre-trained ResNet with FPN to extract high-level features from input images. This approach ensures robustness and adaptability across different datasets.
2.3 Dynamic Weight Adjustment
Dynamic loss functions have been explored in several contexts within machine learning. For example, Dynamic Autoencoders (DynAE) [39] employ adaptive loss functions to balance reconstruction and clustering objectives dynamically during training, enhancing clustering performance by adapting to the data’s changing characteristics. Similarly, DFNet [40] utilizes dynamic loss weights to manage class imbalance and refine segmentation accuracy in semantic segmentation tasks. By calculating weights based on the pixel count of each class within a batch, DFNet mitigates the imbalance and improves performance, particularly for minority classes.
The application of dynamic loss functions in image segmentation, particularly in unsupervised settings, has received relatively less attention. Most existing segmentation methods rely on fixed weights, such as Differentiable Feature Clustering [17], which optimizes pixel labels and feature representations using a combination of feature similarity and spatial continuity constraints. However, achieving the desired segmentation result with this approach often requires manual parameter tuning to find the optimal balancing weight ($\mu$) between feature similarity and spatial continuity. This manual tuning process can be cumbersome and may fail to balance the two constraints well, especially when dealing with images of varying levels of detail and datasets with different characteristics. Moreover, relying solely on fixed weight factors for feature similarity and spatial continuity constraints can lead to the number of predicted clusters converging to one, making it challenging to adapt to different datasets and requiring extensive fine-tuning. Additionally, the lack of adaptability and the need for manual parameter adjustments make integrating this approach with other networks difficult. In contrast, DynaSeg introduces a unique dynamic loss function specifically designed for unsupervised image segmentation. This model employs a Dynamic Weighting Scheme that dynamically adjusts the weighting parameter ($\mu$) during iterations. By incorporating a silhouette score-based phase to guide the optimization process, our model prevents under-segmentation failures and ensures robust performance across diverse datasets. This dynamic approach not only simplifies the integration of our model with other networks but also enhances its adaptability and scalability, making it well-suited for various segmentation tasks and datasets. One of the key benefits of DynaSeg is its ability to autonomously balance competing objectives, thereby enhancing segmentation quality without the need for manual hyperparameter tuning.
@@ 3.2.2 Pre-trained ResNet with Feature Pyramid Network (FPN) Decoder
We incorporate a pre-trained ResNet-18 model with a Feature Pyramid Network (FPN) decoder to enhance our network’s feature extraction capabilities. While pyramid networks have been utilized in segmentation tasks, such as in APCNet [42] and FPN for Land Segmentation [43], our methodology incorporates several novel elements that distinguish DynaSeg from these existing methods. Unlike traditional pyramid networks [42] that construct multi-scale contextual representations using global-guided local affinity, DynaSeg extends this by dynamically adjusting the segmentation process, effectively balancing feature similarity and spatial continuity to produce more nuanced and context-aware segmentation outputs. This dynamic adjustment addresses the limitations of static parameter settings in conventional pyramid networks. Additionally, our approach uniquely integrates a pre-trained ResNet with an FPN decoder specifically tailored for unsupervised segmentation, enhancing feature extraction through robust multi-resolution processing. optimized for unsupervised learning, distinguishing it from other FPN methods [42] [43] .
@@ 3.3.1 Loss Function
Self-supervised Multi-view Clustering [44] 0.316 0.266 0.391 0.339 0.383
Double Clustering with Superpixel Fitting [45] 0.338 0.291 0.385 0.348 0.376
@@ 4.1.2 BSD500
Method Backbone mIoU All
Modified DC [49] - 9.8
IIC [16] ResNet18 6.7
Picie [34] ResNet18 14.4
DenseSiam [35] ResNet18 16.4
DynaSeg - FSF CNN-Based 30.51
DynaSeg - SCF CNN-Based 27.57
DynaSeg - FSF ResNet-18 + FPN 30.07
DynaSeg - SCF ResNet-18 + FPN 30.52
@@ 4.3.3 BSD500 and PASCAL VOC2012 Datasets
The quantitative results on the BSD500 and PASCAL VOC2012 datasets are detailed in Table 1. DynaSeg’s Spatial Continuity Focus (SCF) achieves a remarkable mIOU of 0.396 on PASCAL VOC2012, the highest among all methods compared, demonstrating its effectiveness in managing complex segmentation challenges. This performance underscores DynaSeg’s adept handling of spatial details and continuity, which are crucial for achieving high-quality segmentation results. DynaSeg’s Feature Similarity Focus (FSF) not only consistently outperforms other methods across the entire BSD500 dataset but also specifically surpasses the Graph-based Segmentation method on the BSD500 Fine dataset, showcasing its robustness and effectiveness in maintaining feature homogeneity across diverse segmentation tasks. This highlights FSF’s distinct advantage over both traditional approaches like Graph-based Segmentation and recent innovations such as Differentiable Double Clustering with Edge-Aware Superpixel Fitting (DDCESF) [45], and other state-of-the-art techniques. In addition to quantitative evaluations, we provide a demo [58] of incremental segmentation using DynaSeg on a sample image from the Pascal VOC2012 dataset. This demo, available in the supplemental video, effectively showcases the practical applications of DynaSeg to real-world images, illustrating its capabilities beyond traditional and contemporary unsupervised segmentation methods.
@@
4.4 Computational Efficiency Experiments
To empirically demonstrate the computational efficiency of DynaSeg, we conducted experiments to compare the computational resources required by different segmentation methods. We measured the total parameters, and floating point operations per second (FLOPs) for each method. The results are presented in Table 4.
Method Total Parameters GFLOPs
DINO + HP [54] 39,641,952 164.15
DINO_Vit_base_16 [59] 86,415,592 16.86
IIC [16] 4,521,024 17.94
PiCIE [34] 23,631,424 4.32
DenseSiam [35] 23,741,558 4.38
STEGO [52] 86,614,089 67.42
HSG [50] 27,968,704 7.13
DynaSeg (CNN_based) 193,900 9.75
DynaSeg (ResNet-18) 12,046,272 1.84
Table 4: Comparison of computational efficiency of different methods.
The results demonstrate that DynaSeg with both CNN-based and ResNet-18 backbones maintains competitive computational efficiency, with lower parameter counts and FLOPs compared to some state-of-the-art methods. Notably, the CNN-based version of DynaSeg demonstrates a good balance between computational cost and performance, making it suitable for applications requiring efficient processing. The results demonstrate that DynaSeg with ResNet-18 is highly efficient, requiring significantly fewer parameters and FLOPs compared to other state-of-the-art methods. For instance, DynaSeg (ResNet-18) uses only 1.84 GFLOPs and 12,046,272. This is significantly lower than methods such as DINO + HP [54], which demands 164.15 GFLOPs and 39,641,952 parameters, and STEGO [52], which requires 67.42 GFLOPs and 86,614,089 parameters. The considerable reduction in computational requirements underscores DynaSeg’s capability to deliver effective performance with much lower computational costs, making it particularly suitable for resource-constrained environments. When examining the CNN-based version of DynaSeg, it maintains a competitive edge with only 193,900 parameters and 9.75 GFLOPs. Although methods like PiCIE [34] and DenseSiam [35] show lower GFLOPs at 4.32 and 4.38 respectively, these efficiencies are achieved through the use of pre-trained models, which DynaSeg does not utilize. This distinction highlights the robustness and inherent efficiency of DynaSeg, as it achieves comparable or better performance without the need for extensive pre-training. The CNN-based DynaSeg, therefore, strikes an excellent balance between computational cost and segmentation performance, further validating its practical applicability.
In summary, the analysis reveals that DynaSeg, in both its ResNet-18 and CNN-based configurations, offers substantial computational advantages over other leading methods. The ResNet-18 version stands out for its minimal computational demands, positioning it as an optimal choice for various segmentation tasks where efficiency is paramount. Meanwhile, the CNN-based version provides a balanced approach, achieving efficiency and high performance independently of pre-trained models. These findings collectively affirm DynaSeg’s effectiveness as a highly efficient and robust solution for unsupervised image segmentation, capable of meeting diverse application requirements while maintaining low computational overhead.